OpenStudy (pulsified333):

Assume that the lottery pays $10 on one play out of 100, it pays $2000 on one play out of 7,500, and it pays $12500 on one play out of 150,000. What probability should be assigned to a ticket's not winning anything?

OpenStudy (pulsified333):

@iLoatheMath do you know how to do this?

OpenStudy (anonymous):

I would assume you divide them to find the probability? Hmm..

OpenStudy (pulsified333):

divide what?

OpenStudy (anonymous):

I am probably doing this all wrong...is there multiple choice?

OpenStudy (pulsified333):

no

OpenStudy (anonymous):

Probability = the number of outcomes that are possible for the terms provided divided by the total number of outcomes possible.

OpenStudy (anonymous):

Have you watched khan academy videos possibly? Those usually help me.

OpenStudy (pulsified333):

no I havent

OpenStudy (anonymous):

I haven't got the slightest clue how to solve this..I would assume you divide each individual one by how much $ lottery gets. This is statistics isn't it? I took Calculus!);

OpenStudy (pulsified333):

its finite

OpenStudy (anonymous):

Oh

OpenStudy (kropot72):

The probability of winning a prize is given by: \[\large \frac{1}{100}+\frac{1}{7500}+\frac{1}{150000}=\frac{1521}{150000}\] Therefore the probability of a ticket's not winning anything is: \[\large 1-\frac{1521}{150000}=\frac{148479}{150000}\]